A STUDY OF NbCr AND NbCr^{−} BY ANION PHOTOELECTRON SPECTROSCOPY

\underline{MELISSA A. BAUDHUIN}, PRAVEENKUMAR BOOPALACHANDRAN, SRIJAY S. RAJAN, and DOREEN G. LEOPOLD , Department of Chemistry, University of Minnesota, Minneapolis, MN 55455.

We report the 488 nm photoelectron spectrum of the NbCr^{−} anion. For the ^{2}Δ ground state of neutral NbCr, the short bond length (1.894 Å) and high bond energy (D_{0} 3.0263(6) eV) measured by R2PI spectroscopy^{a} indicate high order multiple bonding. We find that the NbCr^{−} anion has a ^{1}Σ^{+} ground state, in which the ”extra” electron occupies the (4d)δ bonding orbital, giving a 1σ^{2}1π^{4}1δ^{4}2σ^{2} valence electron configuration and a formal bond order of 6. Low-lying excited states of NbCr (assigned as two ^{2}Σ^{+} states) and NbCr^{−} (^{3}Δ) are also observed. The spectra provide the electron affinity of NbCr, energies of the ^{2}Σ^{+} and the ^{3}Δ excited states, vibrational frequencies for the NbCr and NbCr^{−} ground states and for the ^{2}Σ^{+} excited states, and (from Franck-Condon analyses) differences among the bond lengths of the observed states. These results are compared with our previous data for the Group 5/6 congeners NbMo, VCr, and VMo, and with DFT predictions. We also report results for ongoing experiments on the flow tube reactions of the Group V metals Nb and Ta with butadiene, and the vibrationally-resolved photoelectron spectra of some of the organometallic reaction product anions.

aS. M. Sickafoose, J. D. Langenberg, and M. D. Morse, J. Phys. Chem. A. 104, 3521-3527 (2000).